OpenStudy (anonymous):

a triangle has base 15ft and area 60ft^2 what is the height

OpenStudy (mertsj):

A= 1/2 bh Replace b with 15 and A with 60 and solve for h

OpenStudy (anonymous):

wait what

OpenStudy (mertsj):

What is the formula for area of a triangle?

OpenStudy (anonymous):

umm i really dont no 60^2?

OpenStudy (mertsj):

So look in your book and find out.

OpenStudy (anonymous):

i dont have a book this is on paper

OpenStudy (anonymous):

its a packet

OpenStudy (mertsj):

Look at your packet and find the formula for area of a triangle.

OpenStudy (anonymous):

ok theirs only questions

OpenStudy (mertsj):

Ok then. I will tell you again. The formula for the area of a triangle is: A= 1/2 b h

OpenStudy (mertsj):

What do you suppose A stands for?

OpenStudy (anonymous):

1/2bh

OpenStudy (mertsj):

The A stands for area b stands for base h stands for height.

OpenStudy (mertsj):

Your problem says that the area is 60 so replace A in the formula with 60

OpenStudy (mertsj):

You problem says that the base is 15 so replace b in the formula with 15

OpenStudy (mertsj):

Now your formula says 60 = 1/2 ( 15)(h)

OpenStudy (mertsj):

Solve it for h.

OpenStudy (anonymous):

ugg nvm ugg i dont get this